OpenStudy (anonymous):

Which of the following sequences is an arithmetic sequence? A. 5, 20, 80, 320, . . . B. 6, 9, 5, 8, 4, 7, . . . C. 100, 10, 0.1, 0.01, . . . D. –1, 2, 5, 8, . . .

OpenStudy (rituraj):

d; since 2 - (-1) = 3 5 -2 = 3 8 - 5 =3